a paper towel tube has a diameter of 1.7 inches and a height of 11 inches. If the tube were cut and unfolded to form a rectangle, what would be the area of the rectangle?
Answer by nerdybill(7384)

The "width" of the rectangle would be the "circumference" of the tube whereas the length is merely the height of the tube.
.
Therefore, we have:
length = 11 inches
width = 2(pi)r = 2(3.14)(1.7) = 2(3.14)(1.7/2)
width = (3.14)(1.7)
.
area of rectangle = length * width
area of rectangle = 11(3.14)(1.7) = 58.718 square inches